If my sister takes twice as long as I to do her taxes but we can do her taxes together in 10 hours, how long will it take me to do her taxes alone?
I tried to do it and I came up with 2x + x = 10, 3x = 10, x = 3.33
I know this is not correct because it should take me longer than 10 hours and I think the answer is 13.33 hours for me to do her taxes but I don't know how to get the answer. I am missing something.

Answer by stanbon(75887)

If my sister takes twice as long as I to do her taxes but we can do her taxes together in 10 hours, how long will it take me to do her taxes alone?
------------------------
Your DATA:
time = x hr/job ; rate = 1/x job/hr
----------------
Sister DATA:
time = 2x hr/job ; rate = 1/2x job/hr
------------------
Together DATA:
time = 10 hr/job ; rate = 1/10 job/hr.
------------------------
EQUATION:
rate + rate = together rate
1/x + 1/2x = 1/10
Multiply thru by 10x to get:
10 + 5 = x
---------------
x = 15 hrs. (time it will take you alone)
2x = 30 hrs. (time it will take your sister alone)
======================
Cheers,
Stan H.